OpenStudy (anonymous):

how do you write 8 times the quantity d minus 3

OpenStudy (akashdeepdeb):

d minus 3 = d-3 NOW 8 TIMES THAT = 8*(d-3) Understood? :)

OpenStudy (anonymous):

can you help with more